OpenStudy (anonymous):

Which spreadsheet would you use to compute the first nine terms of the geometric sequence:

OpenStudy (anonymous):

\[a_n= 32 (\frac{ 1 }{ 2 }) ^n{-1}\]

OpenStudy (anonymous):

It's supposed to be \[( \frac{ 1 }{ 2 } )^{n-1}\] sorry

OpenStudy (jdoe0001):

\(\large { a_{\color{red}{ n}}=32\left(\frac{1}{2}\right)^{{\color{red}{ n}}-1} \\ \quad \\ \begin{array}{ccllll} term&value\\ \hline\\ 1&a_{\color{red}{ 1}}=32\left(\frac{1}{2}\right)^{{\color{red}{ 1}}-1}\\ 2&a_{\color{red}{ 2}}=32\left(\frac{1}{2}\right)^{{\color{red}{ 2}}-1}\\ 3&a_{\color{red}{ 3}}=32\left(\frac{1}{2}\right)^{{\color{red}{ 3}}-1}\\ ...&... \end{array} } \)